OpenStudy (anonymous):

f(x)=x-2/x^3-2x^2-9x+18 find domain

OpenStudy (phi):

When you see division REMEMBER: NO DIVIDE BY ZERO in this case, where you have x^3-2x^2-9x+18 find all the values of x that make it zero. Solve x^3-2x^2-9x+18=0 this is cubic (highest power is 3) so there are 3 values for x (may be repeated, 2 may be complex) that would make it zero, so are not allowed i.e. not in the domain

OpenStudy (anonymous):

how will I understand this concept? I m not getting any thing?

OpenStudy (phi):

do you understand f(x)= x (a very simple function) for every x, you figure out f(x) (here f(x) is just x) all x will work, so the domain is all real numbers

OpenStudy (anonymous):

ook .

OpenStudy (anonymous):

so all the values with x are function of x

OpenStudy (phi):

a function of x means: given a number (called x) calculate f(x) using that number. Do not over think it. f(x) is defined by a "rule" like f(x)= 2*x+1 or x*x +3, and so on.

OpenStudy (anonymous):

ohhh so fx is the x

OpenStudy (phi):

Now sometimes the rule for f(x) does not work for some x's you know 1/0 is undefined, right? so dividing by 0 is not allowed. if you have a function like f(x)= 1/x this will work for any x except x=0. at x=0 f(x) is undefined. We do not want undefined values, so we exclude x=0. we say the domain is all real numbers except 0

OpenStudy (anonymous):

ok . so f(x)=x-2/x^3-2x^2-9x+18 how we end up with an answer x not equal to -3? and x =2?

OpenStudy (phi):

what is the bottom of that when x=2 ?

OpenStudy (anonymous):

no I mean x not equal to 2 and 3?

OpenStudy (anonymous):

x/xnot equal to 3 x/x not equal to 2?

OpenStudy (phi):

I know what you mean. what is the bottom of that when x=2 ? in other words, what is x^3-2x^2-9x+18 when you replace x with 2?

OpenStudy (anonymous):

then I have to calculate and might be come up with some number

OpenStudy (phi):

yes

OpenStudy (anonymous):

so x has to be 0?

OpenStudy (anonymous):

I donot suppose to come up with any number?

OpenStudy (phi):

what is x^3-2x^2-9x+18 when you replace x with 2?

OpenStudy (anonymous):

(2)^3-2(2)^2-9(2)+18

OpenStudy (phi):

yes, but what number is that?

OpenStudy (anonymous):

8-8-18+18

OpenStudy (anonymous):

0

OpenStudy (phi):

that is the crucial fact. because f(x)=x-2/x^3-2x^2-9x+18 with x=2 gives you \[ f(0)= \frac{0}{0} \] and that is indeterminate (an undefined number). We do not want undefined numbers for f(x). So the way to deal with it is say x=2 NOT ALLOWED same for the other x values (x= 2, x=3, x=-3 not allowed) that make the bottom 0. NO DIVIDE BY ZERO allowed, so don't allow those x values.

OpenStudy (anonymous):

ahan. so x should not be equal to 0 . so to find the domain do we have to first find it with 0, then 1 and then 2 to find out if the answer is 0 or not?

OpenStudy (phi):

so x should not be equal to 0 NOT QUITE You do not what to end up dividing by zero. if you had \[ f(x)= \frac{1}{x-1} \] you do not want x=1 because 1-1=0 and you end up dividing by zero So to answer your question: set the denominator = 0 and solve for all x that make it zero. Those are the x's we do not allow. for your problem: x^3-2x^2-9x+18=0 find all x that make this equation true. Those are the "bad x's" This is a hard problem. But if they give you numbers, you can try them and see if any are "bad"

OpenStudy (anonymous):

so for each question if they donot have given use the values so I have to kep putting values 0, 1 2 3 to find the bad x

OpenStudy (phi):

Guessing might work.... But hopefully the problems they are giving you are "easy" because they are teaching the idea of domain (all "good" x's, no "bad x's") so I hope they don't give hairy problems where it's very difficult to find the "zeros" Post another question....

OpenStudy (anonymous):

so I just guess and write? becuase there is no value of x given

OpenStudy (phi):

which problem?

OpenStudy (anonymous):

this one and others like that?

OpenStudy (anonymous):

becuase how far I can go to keep finding bad and goodx

OpenStudy (phi):

post the next question